# NeuroSwift: A Lightweight Cross-Subject Framework for fMRI Visual Reconstruction of Complex Scenes

### v1

@@ Abstract.
copyright: noneconference: ACM Multimedia Asia Conference; December 9-12, 2025; Kuala Lumpur, Malaysiaisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Human computer interaction (HCI)$c$s: Computing methodologies Artificial intelligence
@@ 1. Introduction
Human brain processes complex visual information efficiently, establishing multi-level neural representations [1, 2, 3, 4]. Early approaches used VAEs or GANs for image synthesis [5, 6, 7, 8, 9, 10] but lacked semantic accuracy. Following the NSD dataset [11], Takagi et al. [12] pioneered high-resolution reconstructions using diffusion models. Subsequent works [13, 14, 15] and recent models [16, 17, 18, 19] employed conditional diffusion and cross-subject decoding to enhance semantic integration and generalizability.
@@
However, existing methods generally suffer from two limitations. First, although models such as MindEye2[16] (64.4M parameters) and MindTuner[17] (76.6M parameters) have achieved cross-subject decoding capabilities, they demand high computational resources. For example, training MindEye2 requires eight NVIDIA A100 GPUs. Nevertheless, MindEye2’s reconstruction performance in cross‑subject generalization still falls short of the results it achieves when trained individually on each subject. Second, in complex visual scenes involving cluttered backgrounds, occlusions, small objects, and dense spatial structures, existing methods often struggle to reconstruct low-level details accurately while also capturing high-level semantic information.
@@ 2.1. Diffusion Models
Diffusion Models (DMs) [20, 21, 22, 23, 7, 14] generate samples through iterative denoising. Latent Diffusion Models (LDMs) operate in compressed latent spaces to reduce computational costs, with Stable Diffusion (SD) [24] being a widely-adopted LDM that leverages large-scale image-text datasets [25] to achieve superior text-to-image performance [26, 8, 27]. We employ SD to transform captions into semantic images. For final reconstruction, we use Versatile Diffusion (VD) [28] which employs a dual-guidance mechanism fusing CLIP features during denoising, trained on Laion2B-en [29] and COYO-700M [30], selected for its ability to integrate multimodal embeddings.
@@
Decoding visual stimuli from brain activity progressed from pre-deep learning methods using image features like multi-scale local image bases[31, 32] and Gabor filters[33] with linear mappings, though limited by fMRI’s low signal-to-noise ratio and small datasets. Deep neural networks later modeled nonlinear relationships [34] through architectures including deep belief networks[35], VAEs[36], feedforward networks[37, 38], GANs[39, 7], and hybrid VAE/GANs[10]. While these improved pixel-level reconstruction, they lacked semantic content until CLIP integration: Lin et al.[5] aligned fMRI patterns with CLIP’s latent space via contrastive-adversarial learning, subsequently using StyleGAN2 to enhance semantic accuracy.
@@
Since Stable Diffusion’s successful application in various generative tasks[40, 25, 41, 20, 42, 32], Takagi et al.[12] pioneered mapping fMRI signals to diffusion latent space and CLIP text embeddings to generate images, though reconstructions lacked sufficient semantic information and natural qualities. Later, MindEye[15], which optimized the semantic representations of fMRI features through contrastive learning, and MindDiffuser[23], which devised a two-stage diffusion process, addressed this issue.
Recent advances in cross-subject decoding include MindEye2[16] enabling cross-subject alignment, MindTuner[17] bridging semantic gaps via visual fingerprinting and fMRI-to-text alignment, and BrainGuard[18] supporting privacy-preserving collaborative decoding. However, these methods still struggle to balance low-level structural feature extraction with high-level semantic understanding, particularly in complex visual scenes under cross-subject generalization, while requiring substantial computational resources.
@@ 3.1. Structural Generation
The AutoKL Adapter transforms fMRI voxels to the latent space $\mathit{z}_{\scriptstyle{\textit{pred}}}$, and the frozen pre-trained AutoKL Decoder decodes it into a predicted image. We call it ”structural generation” because it captures fundamental spatial structures and color distributions that serve as a latent diffusion prior, but it lacks detailed semantic content. See Appendix C for details of the AutoKL Adapter components.
@@ 3.2. Semantic Reinforcement
We use semantic images instead of original visual stimuli for CLIP image alignment because complex natural scenes often contain irrelevant content that may interfere with semantic extraction. In contrast, semantic images generated from textual captions retain only core information, thus reinforcing effectiveness. See Appendix D for details of the CLIP Adapter components.
Training objectives. The CLIP Adapter framework employs two types of loss functions to optimize the mapping of CLIP image and text embeddings. Among them, the SoftCLIP loss[15, 43] has demonstrated effectiveness in aligning the fMRI modality with the pretrained CLIP embedding space. Within this framework, the loss function leverages a contrastive learning mechanism to maximize the similarity of positive pairs while minimizing the similarity of negative pairs.
@@ 4.1. Data Elaboration
We utilize the Natural Scenes Dataset (NSD) [11], a 7T fMRI dataset of eight subjects viewing COCO images [44]. Following established methodologies [6, 14, 13, 15, 12, 27], we analyze four subjects (Subj01, 02, 05, 07) with complete sessions. Our design includes: (a) 982-image shared test set; (b) subject-specific training/validation sets (8,859 images each). Preprocessed fMRI (1.8mm resolution) was masked using NSD-provided ROIs to extract ventral visual cortex voxels. The details of the data used in our experiments have been summarized in Appendix A.2.
Notably, previous studies have primarily relied on standardized brain templates for spatial parcellation. However, both neuroanatomical architecture and functional organization demonstrate significant inter-individual variability[45]. During spatial normalization, these inter-subject discrepancies can induce boundary blurring or registration inaccuracies, ultimately compromising model training efficacy and image reconstruction performance. In contrast to conventional approaches, our study utilizes NSD provided brain masks incorporating manually delineated ROI boundaries for individual subjects. These customized masks account for neuroanatomical and functional specificity through fine grained regional mapping, enabling differentiation of areas with subtle functional distinctions, thereby enhancing downstream processing precision.
@@ 4.3. Qualitative Comparison
In the qualitative experiments, we first compare the reconstruction results of NeuroSwift, fine-tuned with 1 hour of training data under the cross-subject generalization setting, against MindTuner[17] and MindEye2[16] under the same condition. We then compare our single-subject reconstruction results with those of MindBridge[43], trained with 40 hours of data under the single-subject setting.
Figure 4. Comparison of our framework, MindEye2[16], and MindTuner[17] using 1h of training data in cross‑subject adaptation mode.
Figure 5. Comparison of our framework with MindBridge[43] using 40h of training data in single‑subject mode.
@@ 4.4. Quantitative Evaluation
To objectively measure reconstruction performance, we conduct quantitative evaluations benchmarking against four state-of-the-art approaches: MindEye2[16], MindBridge[43], MindTuner[17] and BrainGuard[18]. We assess low‑level spatial coherence with SSIM [46] and pixel-wise accuracy via PixCorr [46], mid‑level texture consistency using AlexNet(2/5) feature similarity [44], and high‑level semantics with Incep [47], CLIP [48], EffNet-B [36], and SwAV [49]. Higher scores are better for SSIM, PixCorr, Alex(2), Alex(5), Incep and CLIP, while lower distances indicate superior performance for EffNet-B and SwAV.
@@ 4.6. Interpretability of NeuroSwift
To investigate the importance of different brain regions in the NeuroSwift for visual reconstruction, we extracted the weight matrices of the AutoKL and CLIP Adapters’ fully connected layers and computed each voxel’s L2 norm as its activation contribution. Figure 7 shows the results obtained by training on 40 hours of data from Subj01 and fine-tuning with 1 hour of data from Subj02. The cross-subject fine-tuning results for Subj05 and Subj07 are provided in Appendix B.
@@ 5. Conclusion and Limitations
Current limitations include dataset-specific validation, as significant differences in experimental design and signal acquisition restrict generalizability to the NSD dataset[11]. Future work will conduct cross-dataset validation on BOLD5000[50] to enhance robustness.
Appendix A Visual mechanisms and ROI selection
A.1. Visual neural mechanism
Our hierarchical visual reconstruction framework inspired by the brain’s multi-level visual processing mechanism. The process of turning light into vision begins in the eye. Light enters through the cornea and projects the retina, where photoreceptors convert it into neural signals. These signals are transported via the optic nerves to subcortical areas in the thalamus, such as the lateral geniculate and pulvinar nuclei, which then send the information mainly to the occipital lobe in the back of the brain. The occipital lobe is divided into several regions, including the primary visual cortex (V1) and interconnected higher-order visual areas. Therefore, we selected these regions as ROIs (see Figure 8).
Figure 8. Graphical representation of the main cortical regions involved in visual perception. The image is cited from Lonta[4].:
A.2. Functions of selected ROIs
We selected several regions of interest (ROIs) to capture key stages of visual processing. Table 4 outlines their principal functions and anatomical locations, which we describe in more detail below.
The Early visual cortices in the occipital lobe (V1, V2, V3) are responsible for processing low-level visual features such as edges, orientation, and contrast. The MidVentral area V4 primarily handles color and shape integration, while the MidLateral region V5/MT specializes in motion perception, encoding speed and direction. Additionally, the posterior parietal cortex (PPC), the MidParietal lobe, plays a crucial role in spatial attention, visuospatial positioning, and integrating multimodal visual information.
Further along the visual pathways, high-level areas contribute to advanced perceptual functions. The inferior temporal cortex (ITC) and fusiform face area (FFA) in the Ventral stream are responsible for object and face recognition, respectively, supporting high-level visual cognition. In the Lateral regions, the extrastriate body area (EBA) processes body shapes and biological motion, while the frontal eye fields (FEF) coordinate eye movements and visual attention, enabling integration between visual perception and motor control. In the Parietal region, the posterior parietal cortex (PPC) plays a key role in spatial attention and visuomotor integration. Together, these regions form a hierarchical network supporting the full spectrum of visual processing.
Table 4. Visual Pathway (ROIs) we used, their primary functions, and corresponding anatomical regions.
Table 5. The details of the NSD data we used in our experiments.
Appendix B Interpretability analysis results on other subjects
Figure 9. Spatial mapping of brain region contributions to AutoKL Adapter and CLIP Adapter on Subj05 and Subj07.:
Figure 9 shows the weight distribution maps of both the AutoKL Adapter and CLIP Adapter for Subj05 and Subj07 under the fine-tuning mode with 1 hour of training data. These maps exhibit activation patterns similar to those of Subj02. For the mapping of the AutoKL Adapter, pronounced activation clusters in ”Early” regions were observed in both subjects, with stronger weight contributions compared to ”Midventral”, ”Midlateral”, and ”Midparietal” regions, where activations remained relatively sparse. The mapping of the CLIP Adapter weights retained the continuous and extended activation pattern across “‘Ventral”, “‘Lateral”, and “‘Parietal” regions, consistent with their role in multimodal semantic integration. While “‘Early”, “‘Midventral”, “‘Midlateral”, and “‘Midparietal” regions showed relatively less involvement, echoing the functional dissociation observed in Subj01 and Subj02.
Appendix C Components of the AutoKL Adapter
The AutoKL Adapter first projects fMRI voxels into a hidden space via linear mapping, followed by sequential processing through LayerNorm, SiLU, and Dropout. These representations are then transformed by residual MLP blocks before being passed to a second linear layer that generates low-resolution feature maps. The maps undergo GroupNorm normalization and upsampling to produce the output $\mathit{z}_{\scriptstyle{\textit{pred}}}$ , ensuring alignment with the latent vector $\mathit{z}_{\scriptstyle{\textit{gt}}}$, which is the variable encoded by the AutoKL Encoder from visual stimulus.
Appendix D Components of the CLIP Adapter
The CLIP Adapter framework uses an embedding network to compress inputs into a hidden space via linear projection, normalization, activation, and dropout, while a residual MLP refines features through cascaded blocks for stable gradients and semantic reinforcement. Subsequently separate linear heads project features into CLIP’s multimodal spaces: the image head to $\mathit{e}_{\scriptstyle{\textit{img\_pred}}}$ (CLIP’s image encoder) and the text head to $\mathit{e}_{\scriptstyle{\textit{txt\_pred}}}$ (CLIP’s text encoder), maintaining distinct semantics while preserving CLIP’s contrastive alignment. This architecture separates semantic enhancement and modality projection, maintaining compatibility with CLIP’s pretrained space while enabling authentic semantic information derived from brain signals.

### v2

@@ Abstract.
copyright: nonejournalyear: 2025copyright: ccconference: ACM Multimedia Asia; December 9–12, 2025; Kuala Lumpur, Malaysiabooktitle: ACM Multimedia Asia (MMAsia ’25), December 9–12, 2025, Kuala Lumpur, Malaysiadoi: 10.1145/3743093.3771019isbn: 979-8-4007-2005-5/2025/12ccs: Human-centered computing Human computer interaction (HCI)$c$s: Computing methodologies Artificial intelligence
@@ 1. Introduction
Human brain processes complex visual information efficiently, establishing multi-level neural representations (brain1, ; brain2, ; brain3, ; retina, ). Early approaches used VAEs or GANs for image synthesis (2-7, ; 3-8, ; 4-9, ; 5-10, ; 6-11, ; 7-12, ) but lacked semantic accuracy. Following the NSD dataset (8-14, ), Takagi et al. (9-1, ) pioneered high-resolution reconstructions using diffusion models. Subsequent works (10-2, ; 11-3-23, ; 12-4, ) and recent models (mindeye2, ; mindtuner, ; BrainGuard, ; Neuropictor, ) employed conditional diffusion and cross-subject decoding to enhance semantic integration and generalizability.
@@
However, existing methods generally suffer from two limitations. First, although models such as MindEye2(mindeye2, ) (64.4M parameters) and MindTuner(mindtuner, ) (76.6M parameters) have achieved cross-subject decoding capabilities, they demand high computational resources. For example, training MindEye2 requires eight NVIDIA A100 GPUs. Nevertheless, MindEye2’s reconstruction performance in cross‑subject generalization still falls short of the results it achieves when trained individually on each subject. Second, in complex visual scenes involving cluttered backgrounds, occlusions, small objects, and dense spatial structures, existing methods often struggle to reconstruct low-level details accurately while also capturing high-level semantic information.
@@ 2.1. Diffusion Models
Diffusion Models (DMs) (3-7, ; mdb8, ; Mdb13, ; 14-6, ; 4-9, ; 11-3-23, ) generate samples through iterative denoising. Latent Diffusion Models (LDMs) operate in compressed latent spaces to reduce computational costs, with Stable Diffusion (SD) (mdb30, ) being a widely-adopted LDM that leverages large-scale image-text datasets (3-28, ) to achieve superior text-to-image performance (mdb26, ; 5-10, ; mdb47, ). We employ SD to transform captions into semantic images. For final reconstruction, we use Versatile Diffusion (VD) (mdb46, ) which employs a dual-guidance mechanism fusing CLIP features during denoising, trained on Laion2B-en (uni31, ) and COYO-700M (uni4, ), selected for its ability to integrate multimodal embeddings.
@@
Decoding visual stimuli from brain activity progressed from pre-deep learning methods using image features like multi-scale local image bases(11-3-08, ; add3, ) and Gabor filters(11-2, ) with linear mappings, though limited by fMRI’s low signal-to-noise ratio and small datasets. Deep neural networks later modeled nonlinear relationships (11-8, ) through architectures including deep belief networks(10-4, ), VAEs(10-5-effenet, ), feedforward networks(10-6, ; 10-7, ), GANs(10-8, ; 4-9, ), and hybrid VAE/GANs(7-12, ). While these improved pixel-level reconstruction, they lacked semantic content until CLIP integration: Lin et al.(2-7, ) aligned fMRI patterns with CLIP’s latent space via contrastive-adversarial learning, subsequently using StyleGAN2 to enhance semantic accuracy.
@@
Since Stable Diffusion’s successful application in various generative tasks(3-27, ; 3-28, ; 3-29, ; 3-7, ; add2, ; add3, ), Takagi et al.(9-1, ) pioneered mapping fMRI signals to diffusion latent space and CLIP text embeddings to generate images, though reconstructions lacked sufficient semantic information and natural qualities. Later, MindEye(12-4, ), which optimized the semantic representations of fMRI features through contrastive learning, and MindDiffuser(14-6, ), which devised a two-stage diffusion process, addressed this issue.
Recent advances in cross-subject decoding include MindEye2(mindeye2, ) enabling cross-subject alignment, MindTuner(mindtuner, ) bridging semantic gaps via visual fingerprinting and fMRI-to-text alignment, and BrainGuard(BrainGuard, ) supporting privacy-preserving collaborative decoding. However, these methods still struggle to balance low-level structural feature extraction with high-level semantic understanding, particularly in complex visual scenes under cross-subject generalization, while requiring substantial computational resources.
@@ 3.1. Structural Generation
The AutoKL Adapter transforms fMRI voxels to the latent space $\mathit{z}_{\scriptstyle{\textit{pred}}}$, and the frozen pre-trained AutoKL Decoder decodes it into a predicted image. We call it ”structural generation” because it captures fundamental spatial structures and color distributions that serve as a latent diffusion prior, but it lacks detailed semantic content. See Appendix LABEL:apen.a for details of the AutoKL Adapter components.
@@ 3.2. Semantic Reinforcement
We use semantic images instead of original visual stimuli for CLIP image alignment because complex natural scenes often contain irrelevant content that may interfere with semantic extraction. In contrast, semantic images generated from textual captions retain only core information, thus reinforcing effectiveness. See Appendix LABEL:apen.b for details of the CLIP Adapter components.
Training objectives. The CLIP Adapter framework employs two types of loss functions to optimize the mapping of CLIP image and text embeddings. Among them, the SoftCLIP loss(12-4, ; 13-5, ) has demonstrated effectiveness in aligning the fMRI modality with the pretrained CLIP embedding space. Within this framework, the loss function leverages a contrastive learning mechanism to maximize the similarity of positive pairs while minimizing the similarity of negative pairs.
@@ 4.1. Data Elaboration
We utilize the Natural Scenes Dataset (NSD) (8-14, ), a 7T fMRI dataset of eight subjects viewing COCO images (imagenet-alex, ). Following established methodologies (3-8, ; 11-3-23, ; 10-2, ; 12-4, ; 9-1, ; mdb47, ), we analyze four subjects (Subj01, 02, 05, 07) with complete sessions. Our design includes: (a) 982-image shared test set; (b) subject-specific training/validation sets (8,859 images each). Preprocessed fMRI (1.8mm resolution) was masked using NSD-provided ROIs to extract ventral visual cortex voxels. The details of the data used in our experiments have been summarized in Appendix LABEL:apen.visual.
Notably, previous studies have primarily relied on standardized brain templates for spatial parcellation. However, both neuroanatomical architecture and functional organization demonstrate significant inter-individual variability(visualdiff, ). During spatial normalization, these inter-subject discrepancies can induce boundary blurring or registration inaccuracies, ultimately compromising model training efficacy and image reconstruction performance. In contrast to conventional approaches, our study utilizes NSD provided brain masks incorporating manually delineated ROI boundaries for individual subjects. These customized masks account for neuroanatomical and functional specificity through fine grained regional mapping, enabling differentiation of areas with subtle functional distinctions, thereby enhancing downstream processing precision.
@@ 4.3. Qualitative Comparison
In the qualitative experiments, we first compare the reconstruction results of NeuroSwift, fine-tuned with 1 hour of training data under the cross-subject generalization setting, against MindTuner(mindtuner, ) and MindEye2(mindeye2, ) under the same condition. We then compare our single-subject reconstruction results with those of MindBridge(13-5, ), trained with 40 hours of data under the single-subject setting.
Figure 4. Comparison of our framework, MindEye2(mindeye2, ), and MindTuner(mindtuner, ) using 1h of training data in cross‑subject adaptation mode.
Figure 5. Comparison of our framework with MindBridge(13-5, ) using 40h of training data in single‑subject mode.
@@ 4.4. Quantitative Evaluation
To objectively measure reconstruction performance, we conduct quantitative evaluations benchmarking against four state-of-the-art approaches: MindEye2(mindeye2, ), MindBridge(13-5, ), MindTuner(mindtuner, ) and BrainGuard(BrainGuard, ). We assess low‑level spatial coherence with SSIM (pixcor-ssim, ) and pixel-wise accuracy via PixCorr (pixcor-ssim, ), mid‑level texture consistency using AlexNet(2/5) feature similarity (imagenet-alex, ), and high‑level semantics with Incep (inception, ), CLIP (uni23-clip, ), EffNet-B (10-5-effenet, ), and SwAV (swav, ). Higher scores are better for SSIM, PixCorr, Alex(2), Alex(5), Incep and CLIP, while lower distances indicate superior performance for EffNet-B and SwAV.
@@ 4.6. Interpretability of NeuroSwift
To investigate the importance of different brain regions in the NeuroSwift for visual reconstruction, we extracted the weight matrices of the AutoKL and CLIP Adapters’ fully connected layers and computed each voxel’s L2 norm as its activation contribution. Figure 7 shows the results obtained by training on 40 hours of data from Subj01 and fine-tuning with 1 hour of data from Subj02. The cross-subject fine-tuning results for Subj05 and Subj07 are provided in Appendix LABEL:apen.weights.
@@ 5. Conclusion and Limitations
Current limitations include dataset-specific validation, as significant differences in experimental design and signal acquisition restrict generalizability to the NSD dataset(8-14, ). Future work will conduct cross-dataset validation on BOLD5000(bold5000, ) to enhance robustness.
Acknowledgements.
This work was supported by the Strategic Priority Research Program of Chinese Academy of Sciences (XDB0930302), and the Guangdong Provincial Key Laboratory of Multimodality Non-Invasive Brain-Computer Interfaces (2024B1212010010).